# Streamlined Federated Unlearning: Unite as One to Be Highly Efficient

### v1

@@ III Preliminary
In this section, we introduce the background knowledge and fundamental concepts related to federated learning and the teacher-student model involved in SFU.
III-A Federated learning
FL is a decentralized machine learning approach where multiple clients collaboratively train a global model without sharing their private data. Federated averaging (FedAvg) as a common aggregation method is used in various scenarios [35]. Each client holds its own local dataset $x_{i}$, and the goal is to learn a global model parameter $M^{*}$ that minimizes the overall loss across all clients.
In each iteration, the server sends the current global model $M_{t}$ to all clients. Each client then performs local training on its dataset using stochastic gradient descent (SGD) for $E$ epochs, producing an updated model $M_{i,t+1}$. The server aggregates the model updates from all clients using a weighted average:
$M_{t+1}=\sum_{i=1}^{I}w_{i}M_{i,t+1},w_{i}=\frac{n_{i}}{\sum_{i=1}^{I}n_{i}}\vspace{5pt}$ (1)
where $w_{i}$ is the weight assigned to each client based on the size of its dataset, with $n_{i}$ being the number of samples in the dataset of the client $i$.
The objective is to minimize the global loss function:
$M^{*}=\arg\min_{M}\frac{1}{I}\sum_{i=1}^{I}w_{i}L_{i}(M)\vspace{5pt}$ (2)
III-B Teacher-student model
The concept of the teacher-student model is derived from knowledge distillation [36, 37, 38]. In SFU, we design a multi-teacher system. Instead of completing just a single step, it utilizes multiple teacher models to guide the student model in achieving both influence removal and performance maintenance simultaneously, thus significantly improving unlearning efficiency. The principle of a general teacher-student model is illustrated below:
III-B1 Step 1: knowledge extraction
In the initial phase of the training process, the knowledge base is extracted from the teacher model. Consider a scenario where the teacher model has an output $q$ and the student model has an output $p$. These outputs are typically represented as probability distributions. The raw output of each model is converted into probabilities using the softmax function:
$q=softmax(Z_{T}/T)$ (3)
$p=softmax(Z_{S}/T)\vspace{5pt}$ (4)
III-B2 Step 2: Knowledge transfer
The student model then needs to receive knowledge from the teacher model, thereby reducing the difference between them. In knowledge distillation, the difference between the teacher model and the student model is usually described by KL divergence (Kullback-Leibler divergence). $KL(q||p)$ represents the loss of information when the true distribution $q$ of the teacher model is approximated by the distribution $p$ of the student model. In the teacher-student model, the goal is to minimize this difference so that the student model increasingly approximates the teacher model. Thus, the training objective can be expressed as:
Minimize $\text{Minimize }\,KL(q\|p)=\sum_{i}q_{i}log(q_{i}/p_{i})$ (5)
where $i$ represents the class index, and $q_{i}$ and $p_{i}$ are the predicted probabilities of the teacher model and the student model for the $i$-th class, respectively.
When multiple teachers are involved, it can be understood that the student model learns different knowledge from multiple different teacher models simultaneously. The training objective in this scenario can be defined as:
Minimize $\text{Minimize }\,KL(q_{1}\|p)+\dots+KL(q_{j}\|p)$ (6)
$=\sum_{i}q_{1,i}log(q_{1,i}/p_{i})+\dots+\sum_{i}q_{j,i}log(q_{j,i}/p_{i})$
where $q_{1,i},\dots,q_{j,i}$ and $p_{i}$ are the predicted probabilities of the $j$ teacher models and the student model for the $i$-th class, respectively.
By minimizing the above loss function, the student model can be made to approach the outputs of multiple teacher models, thus integrating the knowledge from multiple teacher models.
IV Approach
In this section, we first introduce the design philosophy of each teacher model in the multi-teacher system used in SFU. Then, we provide a detailed description of the overall structure of the system, and finally, we present the specific implementation of SFU.
IV-A Design and initialization of teacher models
In SFU, three distinct teacher models have been developed to fulfill specific roles and responsibilities, thereby ensuring the efficacy of the unlearning process and the stability of the model’s performance. The framework of SFU is shown in Figure 1.
@@ IV-A Design and initialization of teacher models
  • Teacher Model 3 is designated as the label-based preservation teacher model ($M_{i}^{L}$), which $i$ denotes the $i$-th client. It consists of one-hot encodings of the true labels of the non-target data on each client, making it unique to each client. $M_{i}^{L}$ complements $M^{P}$. Relying solely on $M^{P}$ is insufficient to eliminate the influence of the target data while preventing a significant decline in model performance. Therefore, we devised $M_{i}^{L}$, which aims to align the predictions of the student model on non-target data as closely as possible to the true labels, thereby further preserving the model’s performance. Since $M_{i}^{L}$ is generated based on local data exclusively, no supplementary training is necessary, thereby avoiding additional computational resource consumption.
IV-B Design of the multi-teacher system
@@
$M_{i}^{U}=M^{O}$ (7)
The local dataset consists of $x_{i}$, which can be expressed as the sum of $x_{i}^{r}$ (the target data to be forgotten) and $x_{i}^{f}$ (the non-target data to be retained).
According to the previous definition, the performance preservation teacher model is designated as $M^{P}$, whereas the forgetting teacher model is designated as $M^{F}$. Since the generation of $M^{F}$ only requires the structure of the global original model $M^{o}$ and not its parameter distribution, we can have:
$M^{P}=random\_initialize(M^{O})$ (8)
In light of the aforementioned, it can be posited that the objective of $M_{i}^{U}$ is to facilitate the unlearning of the target data $x_{i}^{f}$, thereby approximating the predictions of $M^{F}$. This is to reduce the discrepancy between the performance of $M_{i}^{U}$ and $M^{F}$ on the target data. Additionally, $M_{i}^{U}$ strives to maintain the performance on the retained data $x_{i}^{f}$ without any deterioration, aligning its predictions closely with the original state, thereby reducing the performance gap between $M_{i}^{U}$ and $M^{P}$ on the non-target data. To quantify the discrepancy between the two models, we have elected to employ the KL-Divergence metric. Consequently, the objective delineated for SFU is presented below:
Minimize $\text{Minimize }\,KL(M_{i}^{u}(x_{i})\parallel(\alpha\cdot M^{p}(x_{i}^{r})+M^{f}(x_{i}^{f})))\vspace{5pt}$ (9)
where $M_{i}^{u}(x_{i})$ represents the predictive distribution of the unlearned model on the local client data, $M^{f}(x_{i}^{f})$ represents the predictive distribution of the forgetting teacher model on the target forgotten data, and $M^{P}(x_{i}^{r})$ represents the predictive distribution of the performance preservation teacher model’s predictive distribution on the retained data. $\alpha=(len(x^{r}))/(len(x^{f}))$ is a customized hyperparameter to balance the quantitative gap between the target data and the retained data to safeguard the unlearning performance. The detailed analysis of $\alpha$ can be found in section V-B4.
@@
$M_{i}^{L}=one\_hot(y_{labels}(x_{i}^{r}))\vspace{5pt}$ (10)
It represents the $one\_hot$ encoding of the true labels of the retained non-target data on the local client. The objective is to narrow the performance gap between $M_{i}^{U}$ and both $M^{F}$ on $x_{i}^{f}$ and $M_{i}^{L}$, thereby ensuring the continued reliability of the unlearned model on the retained non-target data. Accordingly, the second objective of SFU is presented below:
$\text{Minimize }\,KL(M_{i}^{u}(x_{i}))\|(M_{i}^{t}(x_{i}^{r})+\alpha\cdot M^{r}(x_{i}^{f}))\vspace{5pt}$ (11)
@@
Therefore, the total unlearning goal of Client $i$ is to minimize the loss function:
$Loss$ $=KL(M_{i}^{u}(x_{i}))\|(M^{p}(x_{i}^{r})+\alpha\cdot M^{f}(x_{i}^{f}))$ (12)
$+KL(M_{i}^{u}(x_{i}))\|(M_{i}^{t}(x_{i}^{r})+\alpha\cdot M^{r}(x_{i}^{f}))$
IV-C The realization of SFU
In a FL environment, the server maintains a global model, while each client retains its local data and a local model. This study is based on the assumption that the environment is isomorphic, which implies that all clients possess the same model architecture. Data are not shared between clients, and the server and local models are updated periodically solely through the exchange of model parameters.
In the event that a user issues an unlearning request during the standard FL training process, SFU performs unlearning through the multi-teacher system as shown in Algorithm 1. Firstly, upon receiving the unlearning request, the server sends an unlearning directive to each client and transmits the current global model ($M^{O}$), generating and sharing the performance preservation teacher model ($M^{P}$) with all clients. Subsequently, upon receiving the directive, each client categorizes its local data into the target data to be forgotten ($x_{i}^{f}$) and the remaining non-target data ($x_{i}^{r}$). Based on the received performance preservation teacher model ($M^{P}$) and the classified non-target data, each client generates a forgetting teacher model ($M^{F}$) and a label-based preservation teacher model ($M_{i}^{L}$). Using a multi-teacher system, clients locally initiate the SFU unlearning operation and upload the updated model parameters to the server. The server aggregates the updates from all clients and checks whether the unlearning process is complete. If not, it instructs the clients to repeat the aforementioned steps until the unlearning is finalized. Once the unlearning is completed, the server notifies each client to update their local databases by removing the target data to be forgotten ($x_{i}^{f}$) while retaining the remaining non-target data ($x_{i}^{r}$), ultimately restoring the system to its original federated learning training process.
The unlearning process for SFU can be initiated at any stage of the FL training process and may be continued after the completion of the original FL process. In other words, SFU does not negate the original FL training process and can be integrated into any stage of the FL training process. In this process, the server is not required to access the local data of each client and is not influenced by the distribution of client data, which may be non-iid.
Algorithm 1 Algorithm of SFU
@@ IV-C The realization of SFU
4:     Generate three teacher models: ${M_{P}}\leftarrow{M^{O}}$, $M^{F}\leftarrow random({M^{O}})$ and $M_{i}^{L}\leftarrow one\_hot(y_{labels}(x_{i}^{r}))$.
@@
6:     Update $M_{i}^{U}$ by the multi-teacher system based on Eq.12.
@@
V Experiments and Analysis
@@
In this section, we first introduce the experimental setup. Next, we conduct a detailed evaluation of SFU and SOTA approaches from the perspectives of efficacy, fidelity, and efficiency, assessing the unlearning performance on both image and text datasets. We then analyze the impact of hyperparameters and provide additional analysis on multiple classes unlearning. Finally, we perform an ablation study on SFU. The experimental results and corresponding analyses are presented in each subsection.
@@
TABLE I: Comparison of the unlearning performance of different unlearning methods. Id of target class represents the label of the target data class that needs to be forgotten on each dataset. Acc on all data represents the prediction accuracy of the model before unlearning. Acc on forgotten/retained data represents the prediction accuracy of the model on target forgotten data and remaining data after unlearning, respectively. BD_ASR represents the success rate of the model by backdoor attack after the target forgotten data is implanted with the backdoor.
@@
Retrain FUCDP SFU Retrain FUCDP SFU Retrain FUCDP SFU
@@
0 96.20 90.24 0.02 - 0.04 96.89 - 96.88 0.89 - 0.69
5 96.44 92.36 0.05 - 0.08 96.23 - 96.54 0.46 - 0.83
9 96.17 91.56 0.01 - 0.24 96.44 - 96.14 0.36 - 0.45
TABLE II: Comparison of the efficiency of different unlearning methods. Communication rounds represent the average number of interactions required for unlearning. Time speed-up represents the speed-up ratio in time compared to each method of retraining.
@@
Retrain FUCDP SFU FUCDP SFU
CIFAR-10 312.60 55.40 1.40 7.61$\times$ 152.4$\times$
CIFAR-100 369.44 93.82 2.20 6.32$\times$ 44.04$\times$
Dbpedia 151.48 - 1.40 - 43.45$\times$
V-A Experiments Settings
V-A1 Datasets and Models
To prove the versatility of SFU, we evaluate our proposed method on both image datasets: CIFAR-10, CIFAR-100 [39] and text dataset: DBpedia [40]. The CIFAR-10 dataset consists of 60,000 32x32 color images, organized into 10 classes with 6,000 images per class. The CIFAR-100 dataset is similar to the CIFAR-10 dataset but has 100 classes with 600 images per class. The DBpedia dataset consists of 560,000 training samples and 70,000 testing samples, organized into 14 classes. Each sample includes a class index, title, and content. And we selected different network models for training on each dataset: ResNet18 for CIFAR-10, ResNet44 for CIFAR-100 [41], and a simple lstm [42] for Dbpesia.
V-A2 Training Settings

### v2

@@ IV-A2 Step 2: Knowledge Transfer
IV-B Design and Initialization of Teacher Models
In SFU, three distinct teacher models have been developed to fulfill specific roles and responsibilities, thereby ensuring the efficacy of unlearning and the stability of the model performance. The framework of SFU is shown in Figure 2.
@@ IV-B Design and Initialization of Teacher Models
  • Teacher Model 3 is designated as the label-based preservation teacher model ($M_{i}^{L}$), which $i$ denotes the $i$-th client. It consists of one-hot encodings of the true labels of the non-target data on each client, making it unique to each client. $M_{i}^{L}$ complements $M^{P}$. Relying solely on $M^{P}$ is insufficient to eliminate the influence of the target data while preventing a significant decline in model performance. Therefore, we devise $M_{i}^{L}$, which aims to align the predictions of the student model on non-target data as closely as possible to the true labels, thereby further preserving the model performance. Since $M_{i}^{L}$ is generated based on local data exclusively, no supplementary training is necessary, thereby avoiding additional computational resource consumption.
IV-C Design of the Multi-teacher System
@@
$M_{i}^{U}=M^{O}.$ (7)
The local dataset consists of $x_{i}$, which can be expressed as the sum of $x_{i}^{r}$ (the target data to be forgotten) and $x_{i}^{f}$ (the non-target data to be retained). According to the previous definition, the performance preservation teacher model is designated as $M^{P}$, whereas the forgetting teacher model is designated as $M^{F}$. Since the generation of $M^{F}$ only requires the structure of the global original model $M^{O}$ and not its parameter distribution, we have:
$M^{P}=random\_initialize(M^{O}).$ (8)
In light of the aforementioned, it can be posited that the objective of $M_{i}^{U}$ is to facilitate the unlearning of the target data $x_{i}^{f}$, thereby approximating the predictions of $M^{F}$. This is to reduce the discrepancy between the performance of $M_{i}^{U}$ and $M^{F}$ on the target data. Additionally, $M_{i}^{U}$ strives to maintain the performance on the retained data $x_{i}^{f}$ without any deterioration, aligning its predictions closely with the original state, thereby reducing the performance gap between $M_{i}^{U}$ and $M^{P}$ on the non-target data. To quantify the discrepancy between the two models, we have elected to employ the KL-Divergence metric. Consequently, the objective delineated for SFU is to minimize:
$KL(M_{i}^{U}(x_{i})\parallel(\alpha\cdot M^{P}(x_{i}^{r})+M^{F}(x_{i}^{f}))),\vspace{5pt}$ (9)
where $M_{i}^{u}(x_{i})$ represents the predictive distribution of the unlearned model on the local client data, $M^{F}(x_{i}^{f})$ represents the predictive distribution of the forgetting teacher model on the target forgotten data, and $M^{P}(x_{i}^{r})$ represents the predictive distribution of the performance preservation teacher on the retained data. $\alpha=num(x^{r}))/(num(x^{f})$ is a customized hyperparameter to balance the quantitative gap between the target data and the retained data to safeguard the unlearning performance. The detailed analysis of $\alpha$ can be found in section VI.
@@
$M_{i}^{L}=one\_hot(y(x_{i}^{r})),\vspace{5pt}$ (10)
which represents the $one\_hot$ encoding of the true labels of the retained non-target data on the local client. The objective is to narrow the performance gap between $M_{i}^{U}$ and both $M^{F}$ on $x_{i}^{f}$ and $M_{i}^{L}$, thereby ensuring the continued reliability of the unlearned model on the retained non-target data. Accordingly, the second objective of SFU is to minimize:
$KL(M_{i}^{U}(x_{i}))\|(M_{i}^{L}(x_{i}^{r})+\alpha\cdot M^{F}(x_{i}^{f})),\vspace{5pt}$ (11)
@@
Overall, in order to synthesize the roles of these teacher models and achieve the goal of removing the influence of the forgotten data while maintaining the performance of the model, the total unlearning goal of Client $i$ is summarized as below:
$KL(M_{i}^{U}(x_{i}))\|(M^{P}(x_{i}^{r})+\alpha\cdot M^{F}(x_{i}^{f}))$ (12)
$+KL(M_{i}^{U}(x_{i}))\|(M_{i}^{L}(x_{i}^{r})+\alpha\cdot M^{F}(x_{i}^{f})).$
IV-D Realization of SFU
Algorithm 1 SFU via Multi-Teacher System
@@ IV-D Realization of SFU
4:     Generate three teacher models: ${M^{P}}\leftarrow{M^{O}}$, $M^{F}\leftarrow random({M^{O}})$ and $M_{i}^{L}\leftarrow one\_hot(y_{labels}(x_{i}^{r}))$.
@@
6:     Update $M_{i}^{U}$ by the multi-teacher system based on Equation 12.
@@
In the FL environment, the server maintains a global model, while each client retains its local data and a local model. This study is based on the assumption that the environment is isomorphic, which implies that all clients possess the same model architecture. Data are not shared between clients, and the server and local models are updated periodically solely through the exchange of model parameters.
In the event that a user issues an unlearning request during the standard FL training process, SFU performs unlearning through the multi-teacher system, as shown in Algorithm 1. Firstly, upon receiving the unlearning request, the server sends an unlearning directive to each client and transmits the current global model ($M^{O}$), generating and sharing the performance preservation teacher model ($M^{P}$) with all clients. Subsequently, upon receiving the directive, each client categorizes its local data into the target data to be forgotten ($x_{i}^{f}$) and the remaining non-target data ($x_{i}^{r}$). Based on the received performance preservation teacher model ($M^{P}$) and the classified non-target data, each client generates a forgetting teacher model ($M^{F}$) and a label-based preservation teacher model ($M_{i}^{L}$). Using a multi-teacher system, clients locally initiate the SFU unlearning operation and upload the updated model parameters to the server. The server aggregates the updates from all clients and checks whether the unlearning process is complete. If not, it instructs the clients to repeat the aforementioned steps until the unlearning is finalized. Once the unlearning is completed, the server notifies each client to update their local datasets by removing the target data to be forgotten ($x_{i}^{f}$) while retaining the remaining non-target data ($x_{i}^{r}$), ultimately restoring the system to its original federated learning training process.
V Theoretical Analysis
This section theoretically analyzes how SFU achieves target data influence removal while preserving the performance of the global model on retained data. Besides, we also analyze the compatibility and flexibility of SFU
V-A Theoretical Analysis for Unlearning Effectiveness
The core objective of SFU is to eliminate the influence of target data $x_{i}^{f}$ while maintaining model utility on non-target data $x_{i}^{r}$. To achieve this, the student model $M_{i}^{U}$ is guided by the multi-teacher system through divergence minimization. Specifically, the optimization objective for target data $x_{i}^{f}$ is dominated by the forgetting teacher $M^{F}$, whose predictions approximate a uniform distribution due to random initialization. The divergence term for target data is formulated as:
$\mathcal{D}_{\text{forget}}=\sum_{x_{i}^{f}}M_{i}^{U}(x_{i}^{f})\log\frac{M_{i}^{U}(x_{i}^{f})}{M^{F}(x_{i}^{f})}.$ (13)
Minimizing $\mathcal{D}_{\text{forget}}$ forces the predictions of the student model on $x_{i}^{f}$ to align with $M^{F}(x_{i}^{f})$. Since $M^{F}$ is untrained on $x_{i}^{f}$, its output distribution $q_{F}=M^{F}(x_{i}^{f})$ is approximately uniform, i.e., $q_{F}\propto\frac{1}{K}$ for $K$ classes. This ensures:
$\lim_{\mathcal{D}_{\text{forget}}\to 0}M_{i}^{U}(x_{i}^{f})\to\text{Uniform}(K),$ (14)
effectively erasing the influence of $x_{i}^{f}$. To validate this, consider the gradient of $\mathcal{D}_{\text{forget}}$ with respect to the parameters $\theta$ of the student model:
$\nabla_{\theta}\mathcal{D}_{\text{forget}}=\sum_{x_{i}^{f}}\left(\frac{\partial M_{i}^{U}(x_{i}^{f})}{\partial\theta}\cdot\left(1-\frac{M_{i}^{U}(x_{i}^{f})}{M^{F}(x_{i}^{f})}\right)\right).$ (15)
As optimization progresses, the gradient drives $M_{i}^{U}(x_{i}^{f})$ toward $M^{F}(x_{i}^{f})$, ensuring convergence to an unlearned state.
V-B Theoretical Analysis for Performance Preservation
In addition to removing the influence of target data $x_{i}^{f}$, SFU ensures that the model performance on non-target data $x_{i}^{r}$ remains intact. This is achieved through the joint guidance of two teacher models: the performance preservation teacher $M^{P}$ and the label-based preservation teacher $M_{i}^{L}$.
V-B1 Role of the Performance Preservation Teacher ($M^{P}$)
The performance preservation teacher $M^{P}$ is a direct copy of the original model $M^{O}$. Its purpose is to ensure that the student model $M_{i}^{U}$ retains the knowledge of non-target data $x_{i}^{r}$. The divergence term for this objective is defined as:
$\mathcal{D}_{\text{preserve}}=\sum_{x_{i}^{r}}M_{i}^{U}(x_{i}^{r})\log\frac{M_{i}^{U}(x_{i}^{r})}{M^{P}(x_{i}^{r})}.$ (16)
Minimizing $\mathcal{D}_{\text{preserve}}$ ensures that the predictions of the student model on $x_{i}^{r}$ align with those of the original model, i.e., $M_{i}^{U}(x_{i}^{r})\approx M^{O}(x_{i}^{r})$. To analyze this, consider the gradient of $\mathcal{D}_{\text{preserve}}$ with respect to the student model’s parameters $\theta$:
$\nabla_{\theta}\mathcal{D}_{\text{preserve}}=\sum_{x_{i}^{r}}\left(\frac{\partial M_{i}^{U}(x_{i}^{r})}{\partial\theta}\cdot\left(1-\frac{M_{i}^{U}(x_{i}^{r})}{M^{P}(x_{i}^{r})}\right)\right).$ (17)
As optimization progresses, the gradient drives $M_{i}^{U}(x_{i}^{r})$ toward $M^{P}(x_{i}^{r})$, ensuring that the student model retains the original model performance on non-target data.
V-B2 Role of the Label-based Preservation Teacher ($M_{i}^{L}$)
The label-based preservation teacher $M_{i}^{L}$ provides direct supervision through the ground-truth labels $y(x_{i}^{r})$, encoded as one-hot vectors. The divergence term for this objective is:
$\mathcal{D}_{\text{label}}=\sum_{x_{i}^{r}}M_{i}^{U}(x_{i}^{r})\log\frac{M_{i}^{U}(x_{i}^{r})}{y(x_{i}^{r})}.$ (18)
Minimizing $\mathcal{D}_{\text{label}}$ ensures that the student model’s predictions on $x_{i}^{r}$ align with the true labels, even after unlearning. The gradient of $\mathcal{D}_{\text{label}}$ with respect to $\theta$ is:
$\nabla_{\theta}\mathcal{D}_{\text{label}}=\sum_{x_{i}^{r}}\left(\frac{\partial M_{i}^{U}(x_{i}^{r})}{\partial\theta}\cdot\left(1-\frac{M_{i}^{U}(x_{i}^{r})}{y(x_{i}^{r})}\right)\right).$ (19)
This gradient drives $M_{i}^{U}(x_{i}^{r})$ toward $y(x_{i}^{r})$, ensuring accurate predictions on non-target data.
V-C Compatibility and Flexibility of SFU
V-C1 Non-Disruptive Workflow Integration
The unlearning process in SFU can be initiated at any stage of the FL training process, whether in the early, intermediate, or near-completion stages. After unlearning is completed, SFU seamlessly resumes the original FL training process without requiring additional steps or modifications to the server-side logic. This non-disruptive design ensures that the FL workflow remains uninterrupted, offering significant practical advantages. The global model $M_{g}$ at round $t$ is updated as follows:
  1. a)
    Normal FL Training (Before Unlearning):
    $M_{g}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{t},$ (20)
    where $M_{i}^{t}$ is the local model of client $i$ at round $t$.
  2. b)
    Unlearning Phase:
    During unlearning, the local model $M_{i}^{t}$ is updated using the multi-teacher system for r rounds:
    $M_{i}^{t+r}=M_{i}^{t}-\eta\sum_{k=1}^{r}\nabla_{\theta}\mathcal{D}_{\text{total}}^{(k)}.$ (21)
    where $\mathcal{D}_{\text{total}}$ is the combined divergence term.
  3. c)
    Resuming FL Training (After Unlearning):
    Once unlearning is complete, the global model resumes normal FL training:
    $M_{g}^{t+r+1}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{t+r}.$ (22)
This formulation demonstrates that SFU integrates unlearning into the FL workflow without requiring additional rounds or modifications to the aggregation logic.
V-C2 Privacy Preservation and Robustness
In SFU, the server only receives model updates from clients, which are highly compressed representations of local data. Since the unlearning process is performed locally, no raw data is transmitted to the server. This design preserves data privacy and complies with privacy-centric principles in FL. Moreover, SFU is robust to non-iid (independent and identically distributed) data distributions across clients, as each client performs unlearning independently based on its local data partition. Let $x_{i}=\{x_{i}^{r},x_{i}^{f}\}$ denote the local data partition of client $i$, where $x_{i}^{r}$ is the retained data and $x_{i}^{f}$ is the target data to be forgotten. The unlearning process for client $i$ depends only on $x_{i}$, making SFU invariant to data distribution skewness across clients.
V-C3 Computational and Storage Efficiency
The unlearning process of SFU does not require storing any intermediate information. The three teacher models ($M^{P}$, $M^{F}$, and $M_{i}^{L}$) are generated on-the-fly using lightweight operations: $M^{P}$ is a direct copy of the original model $M^{O}$, $M^{F}$ is initialized randomly using the architecture of $M^{O}$ and $M_{i}^{L}$ is derived from local labels via one-hot encoding. While the introduction of teacher models incurs additional computational overhead during local unlearning, SFU requires only a small number of training rounds to complete the process. Compared to current FU methods, SFU achieves significant efficiency gains.
VI Experiments and Analysis
@@
In this section, we first introduce the relevant settings for the experiments. Then, we conduct detailed experiments for SFU and SOTA approaches from the perspectives of efficacy, fidelity, and efficiency to evaluate the unlearning effect of SFU on both image and text datasets. Additionally, we conduct ablation experiments for combinations of teacher models. Finally, we also experimentally analyze the effect of hyperparameters and supplement the analysis with the effect of unlearning on multiple classes of data. The experimental results and corresponding analyses are presented in each subsection.
@@
TABLE III: Comparison of the unlearning performance of different unlearning methods. Id of target class represents the label of the target data class that needs to be forgotten on each dataset. Acc on all data represents the prediction accuracy of the model before unlearning. Acc on forgotten/retained data represents the prediction accuracy of the model on target forgotten data and remaining data after unlearning, respectively. BD_ASR represents the success rate of the model by backdoor attack after the target forgotten data is implanted with the backdoor.
@@
Retrain FUCDP FUR SFU Retrain FUCDP FUR SFU Retrain FUCDP FUR SFU
@@
0 96.20 90.24 0.02 - - 0.04 96.89 - - 96.88 0.89 - - 0.69
5 96.44 92.36 0.05 - - 0.08 96.23 - - 96.54 0.46 - - 0.83
9 96.17 91.56 0.01 - - 0.24 96.44 - - 96.14 0.36 - - 0.45
TABLE IV: Comparison of the efficiency of different unlearning methods. Communication rounds represent the average number of interactions required for unlearning. Time speed-up represents the speed-up ratio in time compared to each method of retraining.
@@
Retrain FUCDP FUR SFU FUCDP FUR SFU
CIFAR-10 312.60 55.40 18.20 1.40 7.61$\times$ 20.45$\times$ 152.40$\times$
CIFAR-100 369.44 93.82 32.25 2.20 6.32$\times$ 18.62$\times$ 44.04$\times$
Dbpedia 151.48 - - 1.40 - - 43.45$\times$
VI-A Experiments Settings
VI-A1 Datasets and Models
To prove the versatility of SFU, we evaluate our proposed method on both image datasets: CIFAR-10, CIFAR-100 [43] and text dataset: DBpedia [44]. The CIFAR-10 dataset consists of 60,000 32x32 color images, organized into 10 classes with 6,000 images per class. The CIFAR-100 dataset is similar to the CIFAR-10 dataset but has 100 classes with 600 images per class. The DBpedia dataset consists of 560,000 training samples and 70,000 testing samples, organized into 14 classes. And we selected different network models for training on each dataset: ResNet18 for CIFAR-10, ResNet44 for CIFAR-100 [45], and a simple LSTM [46] for Dbpesia.
VI-A2 Training Settings